in a class of 19 mathematicians and 13 chemists two learners are randomly selected one after the other what is the probability that the selected learners are mathematicians?
help me this one pliz.

Answer by Alan3354(69443)

in a class of 19 mathematicians and 13 chemists two learners are randomly selected one after the other what is the probability that the selected learners are mathematicians?
---------
19 + 13 = 32
The 1st pick is 19/32.
The 2nd is 18/31.
---
(19/32)*(18/31) = 171/496